OpenStudy (anonymous):

Evaluate 5a + 3b if a = 7 and b = 2

OpenStudy (idealist10):

5(7)+3(2)=35+6=41 Answer: 41

OpenStudy (zbrandz23):

41

OpenStudy (ericanoel912):

41